OpenStudy (anonymous):

Please, can someone help me!! I need to graph a scatter plot involving the height and arm span of 6 people but I don't have the software to do it and I can't download it either, I don't know why? But anyway, the points are: 66 inches tall arm spam 65, 57 inches tall, arm span 58 inches. 61 inches tall, arm span 61.5, 75 inches tall, arm span 75. 62 inches tall, arm span 61.75, 63 inches tall, arm span 64. I also have to determine the line of best fit. So if you could tell me which two points you used to draw the line of best fit, you would be a real life saver!!

OpenStudy (anonymous):

Can anyone help me??? I really need to get this done but if I can't graph it then I'm kind of screwed!! Unless someone knows a good graphing website that actually works and I don't have to pay for it!!! That would be great to!! Get back to me as soon as possible!!! Please and Thank You!!! :D

OpenStudy (stacey):

Do you have excel?

OpenStudy (stacey):

Excel does a line of best fit. If you do not (or if you want a simpler program), go to http://www.padowan.dk/graph/ and download the program. It is free.

OpenStudy (anonymous):

Yeah, I've tried excel, CNET graph, geogebra and MathGV but nothing works!! :'( Every time its almost done downloading it says theres an error and that this program was unable to be installed??? I don't know why, I think my computer is retarded!! Any other suggestions?? I have to find someway to get this done!!

OpenStudy (anonymous):

Can anyone help me on this?

OpenStudy (stacey):

I think the computer needs to checked for a virus.

OpenStudy (stacey):

Using the height as x and span as y, the CNET graph gave f(x)=0.94148936*x+3.9530142

OpenStudy (anonymous):

Do you think you could do it again but with these points? I know I'm asking a lot but nothing on my computer works. Im taking it down to my uncle who repairs computers to see what he can do for me? And can you tell me what two points make up the line of best fit?? 66,65 57,58 61,61.5 75,75 62,61.75 63,64 55,54 62.5,61.5 68,70.75 61.5,63.25 74,77.75

OpenStudy (stacey):

The program gives f(x)=1.1091566*x-6.3141248

OpenStudy (anonymous):

I also have to answer these questions, Idk if you'd be willing to help with these as well? You don't have to though.. Write the equation of the line passing through those two points using the point-slope formula y - y1 = m(x - x1). Show all of your work. Remember to find the slope of the line first. What does the slope of the line represent within the context of your graph? Using the equation that you found in question 2, approximately how tall is a person whose arm span is 66 inches? According to your line of best fit, what is the arm span of a 74-inch-tall person?

OpenStudy (stacey):

The program does not actually have the line go through two of the specific points in your list.

OpenStudy (anonymous):

Btw I really appreciate you doing this for me. I don't know what I would have done, I cant get my computer fixed until maybe next week and I dont have that much time to wait.

OpenStudy (anonymous):

What do I do then? How do I answer the question then?

OpenStudy (anonymous):

How do I even find the line of best fit? I don't understand when it says its I just have to "eyeball it", what is that suppose to mean?

OpenStudy (stacey):

The line of best fit is a line that comes closest to matching all the points.

OpenStudy (anonymous):

so what are the two points? are they the points the line passes through?

OpenStudy (anonymous):

could you help me with those other questions to?

OpenStudy (anonymous):

is there any way you could send me the graph?

OpenStudy (stacey):

Looking at the graph, I would choose two points on the line- one one each end.

OpenStudy (anonymous):

but there is only one point that falls one the line?

OpenStudy (stacey):

They do not have to be points on your list. Maybe (57, 57) and (76, 78)

OpenStudy (anonymous):

ohhh okay

OpenStudy (anonymous):

hold on one sec

OpenStudy (stacey):

Using the two points, your equation of best fit will be slightly different from the one given by the graph program.

OpenStudy (stacey):

Do you know how to use the two points to find the slope?

OpenStudy (anonymous):

im trying to do the other questions and could you tell me if im right when im done with them.

OpenStudy (anonymous):

m = Y2-Y1/X2-X1 right?

OpenStudy (stacey):

Yes.

OpenStudy (anonymous):

Is the slope 21/19?

OpenStudy (anonymous):

or do I have to simplify it?

OpenStudy (stacey):

21/19 is correct and already simplified.

OpenStudy (anonymous):

okay now I plug it into Y-Y1=M(X-X1) right?

OpenStudy (stacey):

Yes

OpenStudy (anonymous):

awww man Im stuck at Y-57=21/19(x-57) I know you distribute 21/19 to 57 but it doesn't look right to me?

OpenStudy (anonymous):

would you get: Y-57=21/19X-1197/19? I guess im asking a pretty basic question.... when you distribute you multiply??? or divide im drawing a blank.....

OpenStudy (anonymous):

do you multiply*

OpenStudy (stacey):

19 goes into 1197 63 times

OpenStudy (stacey):

So what you have is correct. y - 57 = (21/19)x - 63

OpenStudy (anonymous):

and now I subtract 57 from 63and get Y=21/19X-6 right?

OpenStudy (stacey):

Correct.

OpenStudy (anonymous):

and thats point slope form!! :D

OpenStudy (stacey):

slope-intercept form

OpenStudy (anonymous):

OHHH right!! sorry lol. I have one more, what does this mean? what does the slope of the line represent within the context of your graph?

OpenStudy (stacey):

Y-57=21/19 * (x-57) This was the point-slope form.

OpenStudy (anonymous):

ohhh wow

OpenStudy (stacey):

Remember that the 21 is the change in the y, or the arm span. 19 is the change in the x, or the height.

OpenStudy (anonymous):

I cant believe I missed that one....

OpenStudy (stacey):

So the slope means that as the height increases 19 inches the arm span will tend to increase 21 inches.

OpenStudy (stacey):

They are easy to miss until you really learn them well.

OpenStudy (anonymous):

yeah plus i'm not very good in math so its even harder... wait, was that the answer to what does the slope of the line represent within the context of the graph?

OpenStudy (stacey):

Yes

OpenStudy (stacey):

21 is the change in the y, or the arm span. 19 is the change in the x, or the height. So the slope means that as the height increases 19 inches, the arm span will tend to increase 21 inches.

OpenStudy (anonymous):

okay thank you so much :D

OpenStudy (stacey):

You're welcome.

OpenStudy (anonymous):

Have a good night!

OpenStudy (stacey):

You too.

OpenStudy (anonymous):

Ummm maybe im not done.... sorry to keep going on but I have two other questions: Using the equation that you found in question 2, approximately how tall is a person whose arm span is 66 inches? Would I solve this like this: Y-57=21/19X-66? According to your line of best fit, what is the arm span of a 74-inch-tall person? And for this one would I do: Y-74=21/19X-66?

OpenStudy (anonymous):

how would you solve it? im not sure?

OpenStudy (anonymous):

wait 2 would be Y-74=21/19X-57?

OpenStudy (stacey):

I would use the equation y=21/19 x - 6, but the important part to remember is that y is the arm span, so put 66 in for y.

OpenStudy (anonymous):

so would I change it to 66=21/19x-6? I really suck at algebra... lol

OpenStudy (stacey):

You're doing alright in my opinion.

OpenStudy (anonymous):

haha :)

OpenStudy (stacey):

Now solve for x to get the expected height.

OpenStudy (anonymous):

is it 60?

OpenStudy (anonymous):

would you replace the x with a 0?

OpenStudy (stacey):